Consider the reaction

PCl5 (g) −⇀ PCl3 (g) + Cl2 (g)

Kc = 0.0420

The concentration of the products at equilibrium are [PCl3] = 0.120 M and [Cl2] = 0.190 M.

What is the concentration of the reactant, PCl5, at equilibrium?


[PCl5] = ?



1
Expert's answer
2021-06-22T04:42:13-0400

Equilibrium concentration of PCl3= 0.170 M

Equilibrium concentration of Cl2= 0.190 M

The equilibrium constant for the reaction Kc= 0.0420

The equation relating the equilibrium constant to the equilibrium concentrations of the reactants and products is:


"Kc=\frac{PCl3\times Cl2}{PCl5}"


"0.0420=\frac{0.170\times0.190}{PCl5}"


"PCl5=\frac{0.0420}{0.0323}"


PCl5=1.3000